# A randomized small-block Lanczos method for large-scale null space computations

### v1

@@ 4.2 Partial reorthogonalization
$\Theta=\mathrm{diag}\{\Theta_{1},\dotsc,\Theta_{\ell_{0}}\}\quad\text{and}\quad F=[F_{1}^{\mathsf{T}},\dotsc,F_{\ell_{0}}^{\mathsf{T}}]^{\mathsf{T}},$
@@
(25) $BQ_{i}$ $=Q_{i}\Theta_{i}+Q_{\ell_{0}}F_{i}^{\mathsf{T}}+\Delta_{i}\quad\text{for}\quad 1\leq i\leq\ell_{0},$
$BQ_{\ell_{0}+1}$ $=\sum_{i=1}^{\ell_{0}}Q_{i}F_{i}+Q_{\ell_{0}+1}\Xi_{\ell_{0}+1}+Q_{\ell_{0}+2}\Gamma_{\ell_{0}+1}^{\mathsf{T}}+\Delta_{\ell_{0}+1},$
$BQ_{j}$ $=Q_{j-1}\Gamma_{j-1}+Q_{j}\Xi_{j}+Q_{j+1}\Gamma_{j}^{\mathsf{T}}+\Delta_{j}\quad\text{for}\quad\ell_{0}+2\leq j\leq\ell,$
@@
for $11\leq k\leq\ell+$ 1, where $\mathbf{u}$ denotes the machine precision. Different from full orthogonalization, we allow $Q_{i}$ and $Q_{j}$ to be not (numerically) orthogonal to each other for $i\not=j$. The basic idea of partial reorthogonalization is to monitor this loss of orthogonality by examining $\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ for $1\leq k\leq\ell$. If this value exceeds a certain threshold $\mathbf{u}_{\mathsf{RO}}$, typically chosen as $\mathbf{u}_{\mathsf{RO}}=\sqrt{\mathbf{u}}$, we perform reorthogonalization. Because computing $\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ is computationally expensive, we instead aim to calculate and use an upper bound for this value. Reorthogonalization is not performed if the upper bound is below $\mathbf{u}_{\mathsf{RO}}$, potentially cutting down computational costs while maintaining a decent level of orthogonality.
@@
$\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ $\leq\lVert\Gamma_{\ell}^{-1}\rVert\bigl{\lVert}Q_{k}^{\mathsf{T}}BQ_{\ell}-Q_{k}^{\mathsf{T}}Q_{\ell-1}\Gamma_{\ell-1}-Q_{k}^{\mathsf{T}}Q_{\ell}\Xi_{\ell}-Q_{k}^{\mathsf{T}}\Delta_{\ell}\bigr{\rVert}$
@@
We keep track of the recursive inequalities above using a strictly upper triangular matrix $W$, such that each entry $w_{i,j}$ is an upper bound on $\lVert Q_{i}^{\mathsf{T}}Q_{j}\rVert$ for $11\leq i<j\leq\ell+$ 1. Assuming that columns $1,\ldots,\ell$ have been determined, the entries $1,\ldots,\ell$ of column $\ell+1$ are determined by the recurrences above as follows:
@@ 5 Numerical experiments
In this section, we present some numerical results to support the theoretical results and demonstrate the efficiency of our method. All numerical experiments [^1] Scripts to reproduce numerical results are publicly available at https://github.com/nShao678/Nullspace-code in this section are implemented in Matlab 2022b and executed with an AMD Ryzen 9 6900HX Processor (8 cores, 3.3–4.9 GHz) and 32 GB of RAM.
@@ 5.1 Benchmark examples
[…]
@@
$N_{0}$nodes #connected components
3 168692 122
10 39890 66
@@
Name #rows #columns nnz type nullity
Franz7 10164 1740 40424 Rings 113
Franz8 16728 7176 100368 Rings 1713
Franz9 19588 4164 97508 Rings 619
Franz10 19588 4164 97508 Rings 619
GL7d12 8899 1019 37519 $\mathrm{GL}_{7}(\mathbb{Z})$ 59
GL7d13 47271 8899 356232 $\mathrm{GL}_{7}(\mathbb{Z})$ 961
GL7d14 171375 47271 1831183 $\mathrm{GL}_{7}(\mathbb{Z})$ 7939
@@ 5.2 Single-vector Lanczos
We varied the perturbation level $\epsilon$ from $4^{-4}$ to $4^{-10}$ and ran single-vector Lanczos method for $\ell=1000$ iterations to ensure convergence. Table 4 contains $\lambda_{N}(T_{\ell})$, the $N$th smallest eigenvalue of $T_{\ell}$, and norms $\lVert V^{\mathsf{T}}AV\rVert$ and $\lVert AV\rVert$. As expected, $\lambda_{N}(T_{\ell})$ and $\lVert V^{\mathsf{T}}AV\rVert$ are observed to be of the order $\mathcal{O}(\epsilon)$ and $\mathcal{O}(\epsilon^{2})$, respectively.
@@
Figure 1: Convergence history of single-vector Lanczos with $\epsilon=4^{-10}$ for graph Laplacian: Number of converged Ritz values vs. number of iterations.
@@ 5.3.1 Comparison with null
First, we compare Algorithm 2 with the Matlab built-in function null, which is based on the SVD, for the first six matrices from Table 3. The obtained numerical results are presented in Table 5. For all parameter choices, the computed nullity matches the actual nullity from Table 3. It can be observed that restarting is very effective at limiting the dimension of the Krylov subspace. Although increasing the block size increases the number of matrix-vector multiplications, this increase is, for smaller $d$, outpaced by the higher efficiency of block matrix-vector multiplications. Once the block size becomes too large, both the iteration numbers and execution times start increasing. Based on our experiments, a suitable strategy for choosing the block size is $d=\dim\mathcal{K}/16$, where $\dim\mathcal{K}$ is the maximum dimension of the Krylov subspace before restarting. Algorithm 2 is faster than null for all examples, and the difference becomes more significant as the matrix size increases.
@@ Acknowledgments
The authors thank Yuji Nakatsukasa and Rikhav Shah for helpful discussions on this work.

### v2

@@ 4.2 Partial reorthogonalization
$\Theta=\mathrm{diag}\{\Theta_{1},\dotsc,\Theta_{\ell_{0}}\}\quad\text{and}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F=[F_{1},\dotsc,F_{\ell_{0}}]},$
@@
(25) $BQ_{i}$ $=Q_{i}\Theta_{i}+Q_{\ell_{0}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{i}}+\Delta_{i}\quad\text{for}\quad 1\leq i\leq\ell_{0},$
$BQ_{\ell_{0}+1}$ $=\sum_{i=1}^{\ell_{0}}Q_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{i}^{\mathsf{T}}}+Q_{\ell_{0}+1}\Xi_{\ell_{0}+1}+Q_{\ell_{0}+2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{\ell_{0}+1}}+\Delta_{\ell_{0}+1},$
$BQ_{j}$ $=Q_{j-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{j-1}^{\mathsf{T}}}+Q_{j}\Xi_{j}+Q_{j+1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{j}}+\Delta_{j}\quad\text{for}\quad\ell_{0}+2\leq j\leq\ell,$
@@
for $1\leq k\leq\ell+1$, where $\mathbf{u}$ denotes the machine precision. Different from full orthogonalization, we allow $Q_{i}$ and $Q_{j}$ to be not (numerically) orthogonal to each other for $i\not=j$. The basic idea of partial reorthogonalization is to monitor this loss of orthogonality by examining $\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ for $1\leq k\leq\ell$. If this value exceeds a certain threshold $\mathbf{u}_{\mathsf{RO}}$, typically chosen as $\mathbf{u}_{\mathsf{RO}}=\sqrt{\mathbf{u}}$, we perform reorthogonalization. Because computing $\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ is computationally expensive, we instead aim to calculate and use an upper bound for this value. Reorthogonalization is not performed if the upper bound is below $\mathbf{u}_{\mathsf{RO}}$, potentially cutting down computational costs while maintaining a decent level of orthogonality.
@@
$\lVert Q_{k}^{\mathsf{T}}Q_{\ell+1}\rVert$ $\leq\lVert\Gamma_{\ell}^{-1}\rVert\bigl\lVert Q_{k}^{\mathsf{T}}BQ_{\ell}-Q_{k}^{\mathsf{T}}Q_{\ell-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{\ell-1}^{\mathsf{T}}}-Q_{k}^{\mathsf{T}}Q_{\ell}\Xi_{\ell}-Q_{k}^{\mathsf{T}}\Delta_{\ell}\bigr\rVert$
@@
We keep track of the recursive inequalities above using a strictly upper triangular matrix $W$, such that each entry $w_{i,j}$ is an upper bound on $\lVert Q_{i}^{\mathsf{T}}Q_{j}\rVert$ for $1\leq i<j\leq\ell+1$. Assuming that columns $1,\ldots,\ell$ have been determined, the entries $1,\ldots,\ell$ of column $\ell+1$ are determined by the recurrences above as follows:
@@ 19   Set $F=[O,\Gamma_{\ell_{0}}]V_{Z}$ and $Z_{\ell_{0}+1}=[Z_{\ell_{0}},Q_{\ell_{0}+1}]$;
Now, let us discuss detecting convergence without prior knowledge of the nullity $N$. In the Lanczos method, Ritz values typically converge first to the exterior eigenvalues. An intuitive approach is to terminate the algorithm once an unwanted small Ritz value exceeding $3\epsilon$ has converged. However, this approach is slow if $\lambda_{N+2}-\lambda_{N+1}$ is small. Alternatively, we can leverage the individual convergence property of the small-block Lanczos method. Based on the theoretical results discussed in Remark 3.12 for single-vector Lanczos method, as well as numerical observations for its small-block variants, the convergence speed ss—measured by the number of matrix-vector multiplications required for $d$ newly converged Ritz values—tends to be stable. By estimating the convergence speed during computation, we terminate the algorithm if no additional Ritz values have converged after $s$ matrix-vector multiplications.
@@ 5 Numerical experiments
In this section, we present some numerical results to support the theoretical results and demonstrate the efficiency of our method. Throughout this section, we assume that the nullity $N$ is unknown prior to computation. All numerical experiments [^2] Scripts to reproduce numerical results are publicly available at https://github.com/nShao678/Nullspace-code in this section are implemented in Matlab 2022b and executed with an AMD Ryzen 9 6900HX Processor (8 cores, 3.3–4.9 GHz) and 32 GB of RAM.
@@ 5.1 Benchmark examples
(27) $P_{R}=\mathtt{ichol}(A,\mathtt{struct}\text{(`type',`ict',`droptol',$10^{-3}$,`diagcomp',$0.1$)}).$
@@
$N_{0}$ #nodes #connected components $\underline{\sigma}$
3 168692 122 0.027
10 39890 66 0.041
@@
Name #rows #columns nnz type nullity $\underline{\sigma}^{2}$
Franz7 10164 1740 40424 Rings 113 5.5
Franz8 16728 7176 100368 Rings 1713 2.8
Franz9 19588 4164 97508 Rings 619 8.3
Franz10 19588 4164 97508 Rings 619 8.9
GL7d12 8899 1019 37519 $\mathrm{GL}_{7}(\mathbb{Z})$ 59 3.0
GL7d13 47271 8899 356232 $\mathrm{GL}_{7}(\mathbb{Z})$ 961 1.3
GL7d14 171375 47271 1831183 $\mathrm{GL}_{7}(\mathbb{Z})$ 7939 2.9
@@ 5.2 Single-vector Lanczos
We varied the perturbation level $\epsilon$ from $4^{-4}$ to $4^{-10}$ and ran single-vector Lanczos method for $\ell=1000$ iterations to ensure convergence. We recall that symmetric positive semi-definiteness of the graph Laplacian allows us to use $B=A+\epsilon D$ instead of $B=A^{\mathsf{T}}A+\epsilon D$ in Algorithm 1. As a consequence, the perturbation result in Theorem 2.6 becomes $\lVert V_{N}^{\mathsf{T}}AV_{N}\rVert\leq\epsilon^{2}/(\underline{\sigma}^{2}-\epsilon)$. Table 4 contains $\lambda_{N}(T_{\ell})$, the $N$th smallest eigenvalue of $T_{\ell}$, and norms $\lVert V^{\mathsf{T}}AV\rVert$ and $\lVert AV\rVert$. As expected, $\lambda_{N}(T_{\ell})$ and $\lVert V^{\mathsf{T}}AV\rVert$ are observed to be of the order $\mathcal{O}(\epsilon)$ and $\mathcal{O}(\epsilon^{2})$, respectively.
@@
Figure 1: Convergence history of single-vector Lanczos with $\epsilon=4^{-10}$ for graph Laplacian: Number of converged Ritz values vs. number of matrix-vector multiplications.
@@ 5.3.1 Comparison with null
First, we compare Algorithm 2 with the Matlab built-in function null, which is based on the SVD, for the first six matrices from Table 3. The obtained numerical results are presented in Table 5. For all parameter choices, the computed nullity matches the actual nullity from Table 3, and the approximate basis $V\in\mathbb{R}^{n\times N}$ has almost orthonormal columns in the sense of $\lVert V^{\mathsf{T}}V-I\rVert_{\mathsf{F}}\leq nN\mathbf{u}$, where $\mathbf{u}\approx 2.22\times 10^{-16}$ is the machine precision. It can be observed that restarting is very effective at limiting the dimension of the Krylov subspace. Although increasing the block size increases the number of matrix-vector multiplications, this increase is, for smaller $d$, outpaced by the higher efficiency of block matrix-vector multiplications. Once the block size becomes too large, both the iteration numbers and execution times start increasing. Based on our experiments, a suitable strategy for choosing the block size is $d=\dim\mathcal{K}/16$, where $\dim\mathcal{K}$ is the maximum dimension of the Krylov subspace before restarting. Algorithm 2 is faster than null for all examples, and the difference becomes more significant as the matrix size increases. Although this clearly demonstrates the potential of Algorithm 2, the reductions in execution time relative to null are not as dramatic as one might hope; computing larger-dimensional null spaces remains a challenging problem and further improvements are needed.
@@ Acknowledgments
The authors thank Yuji Nakatsukasa and Rikhav Shah for helpful discussions on this work. They also thank the two anonymous referees, who provided useful comments and critical insights.